OpenStudy (anonymous):

what is y=6x-5 and y=x+9

OpenStudy (anonymous):

two lines

OpenStudy (anonymous):

6x-5=x+9 1) Add 5 both sides 6x=x+14 2)Subtract 1x both sides 5x=14 3) divide 5 x=14/5